The motion of a body is given by the equation dv/dt=4-2v.If the body was at rest at t=0,then find speed of body as a function of time.

Grade:

1 Answers

Arun
4 years ago
 
Dear Suraj
 
dv/dt = 4 – 2v
 
dv/(4 – 2v) = dt
 
 – ½ ln (4-2v) = t + c
 
ln ( 4 – 2v) = – 2t + k
 
4 – 2v = A e-2t
 
2v = 4 – Ae-2t